Find the limit. Enter the exact answer. If the limit does not exist, enter NA. ettex lim x+ex - e-* T
Answer #1

lim X enten ex_e-n. = lim lente) ex (ex-ete- = lim (He-?) ю (1-e->) te-a -2 l-e-a Ito -1 ( 0 lim enten No e - e-n = 1